OpenStudy (anonymous):

All solutions on interval (0, 2pi) sin2x=2sinx

OpenStudy (watchmath):

\(2\sin x\cos x-\sin x=0\) \(\sin x(2\cos x-1)=0\) \(\sin x=0\Rightarrow x=\pi\) \(\cos x=1/2\Rightarrow x=\pi/3,4\pi/3\).

OpenStudy (watchmath):

Sorry I made a mistake

OpenStudy (watchmath):

\(2\sin x\cos x-2\sin x=0\) \(2\sin x(\cos x-1)=0\) \(\sin x=0\Rightarrow x=\pi\) \(\cos x=1\)(no solution on (0,2pi)